Hello, i am doing a project for my school and need to draw 3 rectangles where the area is less than the perimeter. Can you help me?
Answer by Theo(13342)

if the length is 1 or 2, then you can make the width any value greater than or equal to 1 and the area will be less than the perimeter.

if the length is greater than 2, then there is a formula that can help you figure out what the width needs to be.

3 examples are:

length = 1 and width = 5
length = 2 and width = 5
length = 3 and width = 4

in the first case, area = 5 and perimeter = 12
in the second case, area = 10 and perimeter = 14
in the third case, area = 12 and perimeter = 14

area equals length times width.
perimeter equals 2 times length plus 2 times width.

the formula for when length is greater than 2 is:

W < (2L) / (L-2)

for example:

when L is equal to 3, W must be smaller than 6.

in the example above, L was equal to 3 and W was equal to 4 which is smaller than 6 and the area was less than the perimeter.

if L was equal to 3 and W was equal to 7, then area would have been equal to 3*7 = 21 and perimeter would have been equal to 2*3 + 2*7 = 6 + 14 = 20.
in that case the area of 21 would have been greater than the perimeter of 20.